# ABC: Achieving Better Control of Multimodal Embeddings using VLMs

### v1

@@ Abstract
Visual embedding models excel at zero-shot tasks like visual retrieval and classification. However, these models cannot be used for tasks that contain ambiguity or require user instruction. These tasks necessitate a multimodal embedding model, which outputs embeddings that combine visual and natural language input. Existing CLIP-based approaches embed images and text independently, and fuse the result. We find that this results in weak interactions between modalities, and poor user control over the representation. We introduce ABC, an open-source multimodal embedding model that uses a vision-language model backbone to deeply integrate image features with natural language instructions. ABC achieves best-for-size performance on MSCOCO image-to-text retrieval and is the top performing model on classification and VQA tasks in the Massive Multimodal Embedding Benchmark. With a strongly unified vision-language representation, ABC can use natural language to solve subtle and potentially ambiguous visual retrieval problems. To evaluate this capability, we design CtrlBench, a benchmark that requires interleaving textual instructions with image content for correct retrieval. ABC advances the state of multimodal embeddings by offering high-quality representations and flexible natural language control. Our model and datasets are available at our project page.
Machine Learning, ICML
@@ 1 Introduction
Visual embeddings have become a foundational representation in computer vision. Image embedding models have become the state of the art for many zero-shot tasks, including visual retrieval (Chen et al., 2024) and image classification (Yu et al., 2022). Since the release of CLIP (Radford et al., 2021), its dual encoder architecture has remained the state of the art for producing high-quality visual embeddings (Yu et al., 2022; Sun et al., 2023; Chen et al., 2024). However, CLIP only supports separately embedding images or text (Radford et al., 2021). Therefore, complex visual embedding tasks which require additional specification are impossible. For example, a CLIP model cannot distinguish which is the correct answer in Figure 1, as both captions are plausible unless the user provides additional instruction. For such tasks, a multimodal embedding, a representation that interleaves vision and natural language, is essential.
@@
We find that existing approaches suffer from two problems: (1) Vague and repetitive instructions. During training the same instructions are repeated and reused, which results in instruction overfitting (Gudibande et al., 2023). (2) Weak interaction between modalities. Previous works fuse embeddings outputted by CLIP models (Zhang et al., 2024; Wei et al., 2023). This approach prevents deeper interaction between modalities, resulting in superficial use of the instructions (Jiang et al., 2024b).
Figure 1: An example use case of our model; information retrieval on specific aspects of a scene.
Figure 2: An overview of our training regime. We use negative mining to augment our pretraining dataset with almost plausible text negatives for each image query. In our instruction finetuning stage, we craft multiple instructions from each image. We use multiple for captions for same image as negatives, the model must use the natural language instruction to choose the best (positive) text candidate for the query.
@@
To this end, we introduce ABC, a model that uses user instruction to control multimodal embeddings. ABC’s vision-language model (VLM) backbone allows it to integrate natural language instructions when crafting visual embeddings. We find that training our model has two fundamental challenges: (1) Extracting useful contrastive embeddings from a pretrained generative VLM. (2) Designing an instruction fine-tuning method that lets users modify multimodal embeddings using natural language instructions. To train ABC, we adopt a multi-stage training process. In the initial pretraining stage, we use contrastive training with carefully selected negatives to develop a model that generates embeddings, similar to CLIP. In the second stage, we train a lightweight adapter using synthetic natural language instructions that correspond to different aspects of the same image. This training process results in a model that produces powerful and flexible multimodal embeddings.
@@
ABC achieves impressive zero-shot performance in retrieval, classification, and visual question answering (VQA) tasks. In MSCOCO (Lin et al., 2015) image-to-text retrieval, our model outperforms all CLIP models containing at most $8$ billion parameters. Furthermore, our model outperforms all other models on the zero-shot classification and VQA splits of MMEB (Jiang et al., 2024b), a multimodal embedding benchmark spanning 19 tasks. Lastly, we design CtrlBench to measure our model’s ability to use natural language instructions to control retrieval. Using CtrlBench, we show that ABC can accomplish visual retrieval tasks that are fundamentally ambiguous without utilizing natural language instructions.
@@
Our contribution is threefold. (1) ABC: an open-source multimodal embedding model that uses natural language instructions to control visual embeddings. We demonstrate that ABC produces powerful embeddings by benchmarking on zero-shot tasks and robustly utilizes natural language instructions to control embeddings. (2) We provide a decoupled methodology for adapting VLMs into SOTA multimodal embedding models. Previous work integrates instructions during the large-scale contrastive training run. We show that these stages can be decoupled, resulting in a lightweight and adaptable instruction fine-tuning stage that requires only $100$ training steps. (3) Lastly, we introduce CtrlBench, a novel benchmark for measuring instruction-controlled retrieval. CtrlBench requires the model to interleave modalities to retrieve the correct response.
2 Related Works
Visual embeddings. Since Radford et al. (2021) introduced CLIP, various aspects of the original model have been improved. DataComp (Gadre et al., 2023) demonstrated the importance of data filtering during contrastive pretraining to create CLIP models. Training CLIP models requires large-scale pretraining and device parallelization to achieve the throughput and batch sizes needed to produce the best performing models (Cherti et al., 2023b; Sun et al., 2023; Yu et al., 2022). InternVL-G (Chen et al., 2024) is the largest and best performing open-weight CLIP model to date, a $14$ billion parameter model trained to create a vision encoder for the InternVL family of VLMs.
Multimodal embeddings. MagicLens (Zhang et al., 2024) and UniIR (Wei et al., 2023) are multimodal embedding models that fuse CLIP embeddings. By combining multiple CLIP models, they take image-text pairs and align them with image-text pairs using contrastive loss. VLM2Vec (Jiang et al., 2024b) and E5-V (Jiang et al., 2024a) also use VLM backbones to produce multimodal embeddings. VLM2Vec uses contrastive training to align text-image queries to text-image candidates; alternatively, E5-V is trained only by aligning text pairs. Jiang et al. (2024b) introduced Massive Multimodal Embedding Benchmark (MMEB), a benchmark for a variety of multimodal embedding tasks.
@@ 2 Related Works
Extracting Embeddings from LLMs. Several architectural modifications have been proposed for adapting LLMs to produce dense embeddings (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024). LLM2VEC (BehnamGhader et al., 2024) changed the attention mask to be bidirectional, which allows information to flow between all tokens, improving embedding quality. Several methods for extracting dense embedding from LLMs have been proposed, these include picking the last token (Jiang et al., 2024b), mean token pooling (BehnamGhader et al., 2024) or an additional attention layer (Lee et al., 2024). In the text-only domain, the Mistral-7b (Jiang et al., 2023) backboned NV-Embed (Lee et al., 2024) has achieved the top score on MTEB (Muennighoff et al., 2023), a popular text embedding benchmark.
Mined negatives and synthetic instructions. Negative mining is a technique in contrastive training where somewhat relevant candidates are included in each batch (de Souza P. Moreira et al., 2024). This improves the quality of embeddings by ensuring they can differentiate between subtly different candidates. This is often used for training text embedding models, and features prominently in NV-Embed’s training (Lee et al., 2024). MegaPairs (Zhou et al., 2024) uses a VLM to create open-ended instructions to train multimodal retrievers. They use this data synthesis pipeline to train two versions of their MMRet model, one uses CLIP (Radford et al., 2021) as its backbone and the other uses LLava-NeXT (Liu et al., 2024a).
@@ 3 ABC
Figure 2 is an overview of our training regime and model architecture. Our training regime consists of $2$ distinct stages; pretraining and instruction fine-tuning. The pretraining stage uses self-supervised training on image-caption pairs to adapt features used for generative modeling into features for dense embeddings. As our pretraining does not require instructions, it can be easily scaled using any large image-captioning data source (Changpinyo et al., 2021; Schuhmann et al., 2022). In the second stage, we train using queries consisting of images and synthetic text instructions. We also revise the positive caption to correspond to the specific aspect of the image that is relevant to the instruction.
@@ 3.1 Model Design
We require that natural language instructions modify image representations, and vice versa. Therefore, an architecture that supports many attention interactions between modalities is ideal. For this reason, we utilize a VLM as our model backbone. To adapt the VLM to output dense embeddings, we make several architectural changes. Following BehnamGhader et al. (2024), we enable bidirectional attention, allowing all tokens to attend to all other tokens. To create our dense embeddings, we mean pool over tokens in the last hidden layer and project the result using a simple residually connected $MLP$ layer given by equation 1.
Figure 3: A sample from our pretraining dataset. The positive caption (green) is the best caption for the image. The mined negatives (red) are relevant but not the best choice.
@@
Where ${\bm{A}}$ and ${\bm{B}}$ are parameter matrices and $g$ is the element-wise SELU function (Klambauer et al., 2017). To train our backbone we apply LoRA (Hu et al., 2021) adapters on both the vision encoder and LLM modules. We optimize the contrastive loss temperature hyperparameter $\tau$ during pretraining, but freeze it during instruction fine-tuning.
At the beginning of the instruction fine-tuning stage, we fuse the pretrained LoRA weights into the base model and freeze the $MLP$ adapter layer. We then initialize a new lower rank LoRA adapter on our LLM backbone.

### v2

# ABC: Achieving Better Control of Visual Embeddings using VLLMs

@@ Abstract
Visual embedding models excel at zero-shot tasks like visual retrieval and classification. However, these models cannot be used for tasks that contain ambiguity or require user instruction. These tasks necessitate an embedding model which outputs can use a natural language instruction to control the representation of a visual embedding. Existing CLIP-based approaches embed images and text independently, and fuse the result. We find that this results in weak interactions between modalities, and poor user control over the representation. We introduce ABC, an open-source multimodal embedding model that uses a vision-language model backbone to deeply integrate image features with natural language instructions. ABC achieves best-for-size performance on MSCOCO image-to-text retrieval and is the top performing model on classification and VQA tasks in the Massive Multimodal Embedding Benchmark. With a strongly unified vision-language representation, ABC can use natural language to solve subtle and potentially ambiguous visual retrieval problems. To evaluate this capability, we design CtrlBench, a benchmark that requires interleaving textual instructions with image content for correct retrieval. ABC advances the state of visual embeddings, outputting high-quality visual representations with natural language control. Our model and datasets are available at our project page:
@@ 1 Introduction
Figure 1: A task from CtrlBench. Conditioning on the natural language instruction is required to unambiguously retrieve a candidate.
@@
Visual embeddings have become a foundational representation in computer vision. Image embedding models have become the state of the art for many zero-shot tasks, including visual retrieval (Chen et al., 2024) and image classification (Yu et al., 2022). Since the release of CLIP (Radford et al., 2021), its dual encoder architecture has remained the state of the art for producing high-quality visual embeddings (Yu et al., 2022; Sun et al., 2023; Chen et al., 2024). However, CLIP only supports separately embedding images or text (Radford et al., 2021). Therefore, complex visual embedding tasks which require additional specification are impossible. For example, a CLIP model cannot distinguish which is the best answer in Figure˜1. All candidates are plausible unless the user provides additional instruction. However, additional conditioning on a natural language question disambiguates the query, making it apparent that the bottom candidate is the best answer (Stelmakh et al., 2022). Conditioning on natural language gives the user greater control over their visual query. This is especially critical for retrieval augmented generation systems, which must unambiguously retrieve a small number of highly relevant facts from large knowledge bases, to support accurate text generation (Amiraz et al., 2025).
@@
We find that existing approaches suffer from two problems: (1) Reliance on generic or human-annotated instructions (Jiang et al., 2024; Zhang et al., 2025). Unlike the abundant noisy image-text data used to train CLIP models Schuhmann et al. (2022); Changpinyo et al. (2021); Sharma et al. (2018), datasets containing multiple modalities and diverse instruction phrases are significantly more scarce and challenging to curate at scale. This is problematic for embedding models, which require massive datasets for the best representations (Cherti et al., 2023a). Often vague instructions are used repetitively during training, resulting in overfitting of the instruction (Wei et al., 2023; Gudibande et al., 2023). (2) Weak interaction between modalities. Previous works fuse embeddings outputted by CLIP models Zhang et al. (2024); Wei et al. (2023). This approach prevents deeper interaction between modalities, resulting in superficial use of the instructions Jiang et al. (2024).
Figure 2: An overview of ABC’s training regime. First, we mine negatives to augment each batch in our pretraining dataset with almost plausible text candidates. Then we pretrain our model using contrastive loss on independently embedded text-image pairs. Once our model has learned to output high-quality uncontrolled visual embeddings, we introduce instructions. We create multiple instructions and candidate captions for a single image, with each one focusing on different aspects of the image. Lastly, we fine-tune using multiple captions for same image, the positive caption for each query corresponds to the natural language instruction. Therefore, ABC uses the instruction to distinguish the correct candidate for the query.
@@
To this end, we introduce ABC, a model that uses natural language instructions to control visual embeddings. ABC’s Vision Large language Model (VLLM) backbone allows it to integrate natural language instructions when crafting visual embeddings. We find that training our model has two fundamental challenges: (1) Extracting useful contrastive embeddings from a pretrained generative VLLM. (2) Designing an instruction fine-tuning method that lets users modify multimodal embeddings using natural language instructions. To train ABC, we adopt a multi-stage training process. In the initial pretraining stage, we use contrastive training with carefully selected negatives to develop a model that generates embeddings, similar to CLIP. In the instruction fine-tuning stage, we train a lightweight adapter to project queries consisting of images and natural language instructions into the embedding space of the pretrained model. Our synthetic instructions to correspond to different aspects of the same image, therefore, ABC learns to use the instruction for fine-grained control. Our training results in a model that produces powerful and controllable visual embeddings.
@@
ABC achieves impressive zero-shot performance in retrieval, classification, and visual question answering (VQA) tasks. In MSCOCO (Lin et al., 2015) image-to-text retrieval, our model outperforms all CLIP models containing at most $8$ billion parameters. Furthermore, our model outperforms all other models on the zero-shot classification and VQA splits of MMEB (Jiang et al., 2024), a multimodal embedding benchmark spanning 19 tasks. Lastly, we design CtrlBench to measure our model’s ability to use natural language instructions to control retrieval. Using CtrlBench, we show that ABC can accomplish visual retrieval tasks that are fundamentally ambiguous without utilizing natural language instructions.
@@
Our contribution is threefold. (1) ABC: an open-source multimodal embedding model that uses natural language instructions to control visual embeddings. We demonstrate that ABC produces powerful embeddings by benchmarking on zero-shot tasks and robustly utilizes natural language instructions to control embeddings. We study how key design choices—including vision encoder resolution, contrastive loss temperature, and VLLM backbone—impact embedding quality. (2) We provide a decoupled methodology for adapting VLLMs into dense embedding models. Previous work integrates instructions during the large-scale contrastive training run. We show that these stages can be decoupled, resulting in a lightweight and adaptable fine-tuning stage that requires only $100$ training steps using synthetic instructions. (3) Lastly, we introduce CtrlBench, a novel benchmark for measuring instruction-controlled retrieval. CtrlBench requires the model to interleave modalities to retrieve the correct response.
2 Background
Visual embeddings. Radford et al. (2021) proposed CLIP, which projects images and text to a shared dense embedding space. Due to its contrastive training, the similarity of vectors is computed as simple dot product (Hadsell et al., 2006). CLIP embeddings are used for tasks such as zero-shot image classification and image-to-text retrieval, by computing the similarity of an image query $\bm{q}$} with a set of candidates ${\bm{c}}\in\mathcal{C}$. An embedding model $f$ maps queries and candidates to vectors in $\mathbb{R}^{n}$. Given a query $\bm{q}$} and a set of candidates $\mathcal{C}\subseteq\mathbb{R}^{n}$ containing the correct answer $\bm{c}$$^{+}$, the retrieved candidate is:
${\bm{c}}^{*}=\arg\max_{{\bm{c}}\in\mathcal{C}}\;f({\bm{q}})\cdot f({\bm{c}})$
The model is succeeds if ${\bm{c}}^{*}={\bm{c}}^{+}$. Alternatively, when measuring metrics such as recall@K, K of the top scoring candidates are returned. Subsequent works have identified several key factors for training training high-quality CLIP models: DataComp (Gadre et al., 2023) demonstrated the importance of data choice during to create high-quality embeddings. (Cherti et al., 2023b) showed that large-scale pretraining and batch size are crucial for producing the best performing models. Lastly, Align (Jia et al., 2021) found that optimizing the setting of the temperature $\tau$, used to scale the contrastive loss function, was crucial.
Multimodal embeddings for Visual Question Answering. Visual embeddings, especially those of diverse scenes containing many objects, offer limited control over the embedding (Wei et al., 2023). However, a multimodal embedding ${\bm{q}}=(q_{i},q_{t})$, jointly conditioned on image $q_{i}$ and text instruction $q_{t}$ allows for flexible user natural language control over embeddings. This allows the user to create embeddings corresponding to specific objects or concepts in an image (Jiang et al., 2024). For a VQA task, the model embeds $(q_{i},q_{t})$ jointly into a dense vector. The model uses cosine similarity to retrieve the maximally similar text candidate $c^{*}\in\mathcal{C}$. As in the unconditioned setting, the model is successful if $c^{*}$ the correct answer. We use the VQA split of MMEB (Jiang et al., 2024), which contains 10 tasks that span text-conditioned classification and retrieval. We also introduce CtrlBench, a benchmark designed to measure text-conditioned visual retrieval when image-only retrieval is ambiguous.
Ambiguity in visual retrieval. Ambiguity is a well studied problem in information retrieval and natural language processing (Keyvan & Huang, 2022). Often, users provide queries that have multiple distinct, nevertheless plausible answers. Several rectifying approaches have been proposed, including using follow-up questions to disambiguate queries (Kim et al., 2023; Aliannejadi et al., 2019) and query refinement (Stelmakh et al., 2022). Following Min et al. (2020), we define an image query $q_{i}$ as ambiguous if its corresponding candidate pool $\mathcal{C}$ contains multiple semantically distinct plausible answers. Additional conditioning, such as on a natural language instruction $q_{t}$, disambiguates $q_{i}$ if the joint query ${\bm{q}}=(q_{i},q_{t})$ has exactly one plausible answer in $\mathcal{C}$ (Stelmakh et al., 2022).
@@ 2 Background
Extracting Embeddings from LLMs. Generative large language models are a source of pretrained bases training dense embeddings models. NV-Embed (Lee et al., 2024) used contrastive training on an LLM backbone to achieve the best performance on the MMEB text embedding benchmark (Muennighoff et al., 2023). Several architectural modifications have been proposed for adapting LLMs to produce dense embeddings Wang et al. (2024a); BehnamGhader et al. (2024); Lee et al. (2024). LLM2VEC BehnamGhader et al. (2024) changed the attention mask to be bidirectional to allow information flow between all tokens. Furthermore, several methods for extracting dense embedding from LLMs hiddens have been proposed. These include picking the last token (Jiang et al., 2024), mean token pooling (BehnamGhader et al., 2024) or an additional attention layer (Lee et al., 2024).
@@ 3 ABC
Figure˜2 is an overview of our training regime and model architecture. Our training regime consists of $2$ distinct stages; pretraining and instruction fine-tuning. The pretraining stage trains on image-caption pairs to adapt features used for generative modeling into features for dense embeddings. As our pretraining does not require instructions, it can be easily scaled using any large image-captioning data source (Changpinyo et al., 2021; Schuhmann et al., 2022). In the second stage, we train an adapter to allow ABC to embed images conditioned on a text instruction, enabling natural language control over visual representations. Using contrastive learning with synthetic instructions, the ABC learns to represent different visual aspects of the same image based solely on textual commands.
@@ 3.1 Model Design
Pretraining Architecture. During pretraining images and text are embedded independently, and aligned using contrastive loss, much like CLIP (Radford et al., 2021). Text and images are converted to hiddens ${\bm{h}}^{(0)}\in\mathbb{R}^{s\times d}$ using the embedding layer and vision encoder, respectively. The LLM module is shared between both modalities. To adapt the LLM to output dense embeddings, we make several architectural changes. Following BehnamGhader et al. (2024), we enable bidirectional attention, allowing all tokens to attend to all other tokens. To create our dense embeddings, we average over the sequence dimension in the last hidden layer ${\bm{h}}^{(l)}$ (equation 1), and project the result ${\bm{x}}\in\mathbb{R}^{d}$ using a residually connected $MLP$ layer (equation 2).
${\bm{x}}$ $=\frac{1}{s}\sum_{i}^{s}{\bm{h}}^{(l)}_{i}$ (1)
@@
Where $\bm{A}$} and $\bm{B}$} are parameter matrices and $g$ is the element-wise SELU function (Klambauer et al., 2017). To train our backbone, we apply LoRA (Hu et al., 2021) adapters on both the vision encoder and LLM modules. We optimize the contrastive loss temperature hyperparameter $\tau$ during pretraining.
Instruction Fine-tuning Architecture. After pretraining, ABC can project queries consisting of images or text into a shared embedding space. However, to control the representation of visual embeddings, we must project queries consisting of images and text instructions into the embedding space. To accomplish this, we train a low-rank LoRA adapter on the LLM component. The LoRA hooks of the instruction adaptor are attached to the MLP and attention layers in each decoder block of the LLM backbone, we freeze all other weights during instruction fine-tuning. The adapter contrastively aligns image and instruction queries with text candidates embedded using the pretrained model. For queries containing both an image and text instruction, we adopt the instruction template from our VLLM backbone (Wang et al., 2024b):
<vision_start><image_tokens><vision_end><im_start>Instruction: {text}<im_end> (3)
We replace <image_tokens> with tokens output by vision encoder. <vision_start> and <vision_end> are special vocabulary tokens used to denote the start and end of the sequence of image tokens. Similarly, <im_start> and <im_end> are used to denote the start and end of text tokens. If no natural language instruction is used, the last token is <vision_end>. We do not back-propagate through text candidate embeddings, only the queries containing an image and instruction. This ensures that our image-text queries share the same features as text and images embedded using the pretrained model. We can easily alternate between embedding with or without instructions by disabling the instruction fine-tuned LoRA adapter. We freeze the temperature $\tau$ during fine-tuning.
3.2 Pretraining Data
Figure 3: The positive caption (green) is the best caption for the image. The mined negatives (red) are relevant but not the best choice.
To create our pretraining dataset we employ negative mining on Conceptual Captions (Sharma et al., 2018). We derive the mined negatives for our dataset as follows: (1) We do a small pretraining run using only in-batch negatives. (2) We use the resulting model to calculate similarity scores between all images and captions in our pretraining dataset. This approach avoids a circular dependence on a third-party embedding model to train our embedding model. Therefore, it is easily extensible to modalities where an existing embedding model is not publicly available. (3) To prevent our negatives from being too similar to our positive samples, we set a similarity threshold $\epsilon\in[0,1]$. We only sample negatives that have a similarity score of at most $\epsilon$ times the similarity score of the correct candidate. We randomly choose our mined hard negatives from the $100$ candidate captions below the threshold. This results in the text candidates shown in Figure˜3. The mined text negatives are clearly relevant, but the correct caption is still the best answer.
In our pretraining run, each batch consists of $N$ image queries (without instructions) and $M$ text candidates. We include MNM-N mined text negatives in each batch. Therefore, each image query has $\frac{M}{N}-1$ corresponding mined negatives. Our pretraining loss function is given by Equation˜4.
$\mathcal{L}_{pretrain}=-\sum_{i}^{N}\log\frac{\exp(f({\bm{q}}_{i})\cdot f({\bm{c}}_{i}^{+})/\tau)}{\sum_{j=1}^{N}(\exp(\frac{f({\bm{q}}_{i})\cdot f({\bm{c}}_{j}^{+})}{\tau})+\sum_{k=1}^{\frac{M}{N}-1}\exp(\frac{f({\bm{q}}_{i})\cdot f({\bm{n}}_{j}^{k})}{\tau}))}$ (4)
For a given query ${\bm{q}}_{i}$, ${\bm{n}}_{i}^{k}$ is the $k_{th}$ mined hard negative candidate and ${\bm{c}}_{i}^{+}$ is its positive candidate. $\tau$ is the temperature hyperparameter used to scale the loss. ${\bm{n}}_{j}^{k}$ is a mined negative for ${\bm{x}}_{i}$ when $i=j$, otherwise it acts as a regular in-batch negative. The embeddings are unit normalized before loss is computed. We scale the loss by $\frac{1}{N}$, the number of image queries in the batch.
3.3 Instruction Fine-tuning Data
To create our instruction fine-tuning dataset we use Visual Genome (Krishna et al., 2016), a dataset comprised of images with captioned bounding boxes. When choosing which bounding boxes to use for each image, we filter by the size of the bounding box (width $\times$ height). We exclude the $5$ largest bounding boxes, as we find they often do not represent specific objects or aspects of the scene. For each image, we randomly choose $4$ bounding boxes and their respective captions. We then prompt GPT-4o (OpenAI, 2024) to create instructions corresponding to each bounding box caption. We sample multiple captions from each image so that they can be used as negatives for each other during instruction fine-tuning. This ensures that each image query has multiple realistic captions in its batch. Therefore, utilizing the instruction is required to choose the correct text candidate unambiguously. We provide the prompt and generation settings used to create our instruction fine-tuning dataset in Appendix˜A. We instruction fine-tune using exclusively in-batch negatives. However, we group all queries that contain the same image into the same batch. This ensures that queries always have multiple relevant text candidates.
4 Experiments